Can you please help me solve this question: the perimeter of a rectangle equals 114 feet. the length equals 3 less than twice the width. Find the demensions of the rectangle.
I have tried guess and check and comparing it to a square. Thank you!

Answer by ankor@dixie-net.com(22740)

the perimeter of a rectangle equals 114 feet. the length equals 3 less than twice the width.
Find the dimensions of the rectangle.
:
We know:
2L + 2W = 114
:
"the length equals 3 less than twice the width."
L = 2W-3
:
Back to the perimeter equation, replace L with (2W-3)
2(2W-3) + 2W = 114
4W - 6 + 2W
4W + 2W = 114 + 6
6W = 120
W = 120/6
W = 20 is the width
:
You should be able to find the length (L)